# Implicit Actor Critic Coupling via a Supervised Learning Framework for RLVR

### v1

@@ 5 Conclusion
In this work, we propose PACS, a novel RLVR algorithm that achieves implicit actor-critic coupling via supervised learning. By treating outcome rewards as supervised targets and parameterizing the score function with policy log probabilities, PACS addresses the sparse reward and training instability challenges in existing RL-based RLVR methods. Extensive experiments on four mathematical benchmark datasets demonstrate that PACS significantly outperforms strong baselines including PPO and GRPO, achieving superior task performance while maintaining healthier policy entropy for sustained exploration. These findings suggest that supervised learning-based implicit actor-critic coupling represents a promising paradigm for advancing RLVR.
Appendix A Baselines
A.1 Proximal Policy Optimization Algorithms (PPO)
@@
$\mathcal{J}_{\text{PPO}}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(\cdot|q)}\frac{1}{|o|}\sum_{t=1}^{|o|}\left\{\min\left[\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}A_{t},\text{clip}\left(\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})},1-\varepsilon,1+\varepsilon\right)A_{t}\right]\right\}$
@@
A.2 Group Relative Policy Optimization (GRPO)
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}$
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}A_{i,t},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)A_{i,t}\right]-\beta D_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\right\}$
@@
Appendix B Hyperparameters Settings
@@ B.2 Training Hyperparameters
Max prompt length 512
Max response length 4,096
@@
Micro batch size per GPU 16 for Qwen2.5-3B, 8 for Qwen2.5-7B
@@
Total training steps 140 for Qwen2.5-3B, 300 for Qwen2.5-7B
@@ B.3 Inference Hyperparameters
Tensor parallel size 4
@@
Appendix C Detailed Experimental Results
This section presents the comprehensive test results of each model on AMC 23, AIME 2024, and AIME 2025, along with corresponding analyses.
C.1 AMC23
The experimental results clearly demonstrate the superior performance of the PACS algorithm on the AMC23. As shown in Table 5 and Table 6, PACS achieves optimal or near-optimal performance across the majority of $\text{pass@}k$ metrics. The advantages of PACS are particularly pronounced at lower $k$ values; for instance, on the Qwen2.5-3B and Qwen2.5-7B models, its pass@1 scores reach 39.63 and 57.83 respectively, surpassing all baselines. This indicates that the proposed algorithm can more effectively enhance the model’s capability to solve problems in the first attempt or within a few trials, which holds significant value in practical applications. As the model scale expands from 3B to 7B, the PACS algorithm demonstrates excellent scalability, with its performance advantages becoming increasingly pronounced. On the Qwen2.5-7B model, PACS not only achieves leading performance in pass@1 but also comprehensively outperforms other baselines across the broad range from $k=1$ to $k=64$, showing substantial performance improvements. For example, at pass@8, PACS (78.17) outperforms the second-best GRPO (72.20) by nearly 6 points.
AMC23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 30.20 42.56 55.10 66.16 75.62 83.88 90.42 94.85 97.10
PPO 38.57 48.65 57.97 66.07 73.25 80.19 86.81 92.09 95.53
GRPO 37.01 46.85 58.14 68.71 76.69 82.61 87.72 92.39 96.44
PACS 39.63 50.56 60.30 68.23 75.97 82.45 88.23 92.70 96.18
@@ C.1 AMC23
Table 5: Results of Qwen2.5-3B on AMC23 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
AMC23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 35.72 49.47 61.05 70.21 78.35 85.53 91.19 95.89 99.18
PPO 51.22 60.01 67.22 72.80 77.52 81.92 85.71 88.67 90.92
GRPO 57.22 63.36 68.21 72.20 76.06 80.72 86.14 90.69 93.34
PACS 57.83 65.57 71.94 78.17 83.83 88.22 91.95 95.51 98.51
    - w/o weight 54.51 62.95 69.65 75.15 79.77 83.77 86.89 88.97 90.62
@@
Table 6: Results of Qwen2.5-7B on AMC23 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
C.2 AIME 2024
On the more challenging AIME 2024 dataset, PACS again demonstrates its effectiveness. On the Qwen2.5-3B model scale, different algorithms show very similar performance with intense competition, yet PACS still achieves leading results in critical intervals. Throughout the entire $\text{pass@}k$ curve, PACS consistently remains in the top tier. On the Qwen2.5-7B model, PACS achieves comprehensive leadership from $k=1$ to $k=128$, obtaining optimal results at nearly all evaluation points. Similar to the AMC experiments, the base model again achieves the highest score at $k=256$, which further confirms the inference that unaligned base models retain higher answer diversity and have the potential to “enumerate” correct answers through massive sampling, though this approach is highly inefficient for practical applications.
AIME 2024 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 4.92 7.90 11.43 15.66 21.11 27.21 33.45 40.95 49.68
PPO 7.55 10.83 14.96 19.70 24.94 30.27 35.69 42.08 50.81
GRPO 8.14 11.45 14.79 18.13 21.74 26.06 31.65 37.75 42.76
PACS 7.98 12.28 16.98 21.49 25.58 30.10 35.87 42.25 48.32
@@ C.2 AIME 2024
Table 7: Results of Qwen2.5-3B on AIME 2024 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
AIME 2024 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 7.26 11.66 17.17 23.46 29.94 36.41 43.77 52.66 62.55
PPO 15.56 19.45 23.51 28.50 34.16 38.71 41.50 43.85 46.46
GRPO 14.10 16.92 20.71 25.37 30.34 34.78 38.50 41.66 45.42
PACS 15.80 19.93 24.39 29.75 36.29 43.90 51.21 56.54 59.78
    - w/o weight 15.72 19.60 22.24 24.57 27.88 32.58 38.77 46.16 53.42
Table 8: Results of Qwen2.5-7B on AIME 2024 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
C.3 AIME 2025
To comprehensively evaluate the performance stability and generalization capability of PACS, we conduct assessments on the latest AIME 2025 benchmark. This dataset represents the current highest difficulty standard for mathematical reasoning tasks, enabling more rigorous examination of algorithm performance under extremely challenging scenarios. On 3B-scale models, PACS demonstrates impressive comprehensive advantages, achieving optimal performance across nearly all evaluation metrics. PACS obtains superior results throughout the entire evaluation spectrum from pass@1 to pass@256.
On 7B models, PACS achieves complete dominance, from pass@1 of 13.33 to pass@256 of 58.22, with every metric substantially outperforming all baseline methods. It is worth emphasizing that the performance gap between PACS and the second-ranked method is highly significant across multiple key metrics. For instance, at pass@16, PACS (36.00) surpasses GRPO (27.57) by over 8 percentage points; at pass@128, the gap reached 12 percentage points (54.48 vs 42.33).
AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 1.73 3.32 6.14 10.62 16.67 23.45 30.69 39.09 47.38
PPO 1.40 2.71 5.07 8.99 14.56 21.11 28.07 35.52 43.51
GRPO 2.62 3.92 5.29 7.07 9.87 14.07 20.18 28.37 36.58
PACS 2.94 4.92 7.63 11.32 16.47 22.91 30.50 39.07 47.18
Table 9: Results of Qwen2.5-3B on AIME 2025 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 3.35 6.12 10.40 16.04 22.32 28.56 34.99 42.07 50.11
PPO 9.61 14.42 19.69 24.89 29.21 32.32 34.94 38.27 42.91
GRPO 8.36 11.82 16.34 21.63 27.57 33.26 38.12 42.33 46.45
PACS 13.33 18.19 23.73 29.65 36.00 42.88 49.45 54.48 58.22
    - w/o weight 10.44 15.12 19.36 22.86 26.48 31.03 36.59 42.39 47.67
Table 10: Results of Qwen2.5-7B on AIME 2025 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
Synthesizing results across the four benchmarks—MATH 500, AMC23, AIME 2024, and AIME 2025—PACS exhibits exceptional cross-task generalization capability and performance consistency. Whether on the relatively straightforward MATH 500 and AMC23 or the extremely challenging AIME 2025, PACS consistently outperforms existing mainstream methods. This conclusively demonstrates that our proposed algorithm possesses robust adaptability and resilience, enabling outstanding performance across mathematical reasoning tasks of varying difficulty levels.
C.4 Ablation Study

### v2

@@ 7 Conclusion
In this work, we propose PACS, a novel RLVR algorithm that achieves implicit actor-critic coupling via a supervised learning framework. By treating outcome rewards as supervised targets and parameterizing the score function with policy log probabilities, PACS mitigates the sparse reward and training instability challenges which are inherent to existing RL-based methods. Extensive experiments on various benchmarks demonstrate that PACS significantly outperforms strong baselines including PPO and GRPO, achieving superior task performance while maintaining healthier policy entropy for sustained exploration. These results highlight PACS as a promising approach for advancing RLVR.
Appendix A Proofs and Theoretical Discussions
In this section, we provide a rigorous theoretical analysis of the PACS framework. We first derive the exact gradient update of the proposed objective function, demonstrating that it inherently recovers an advantage-weighted policy gradient. Subsequently, we analyze the global optimality conditions to explain why PACS maintains solution diversity and resists entropy collapse, a property achieved through implicit regularization rather than explicit penalty terms.
A.1 Derivation of the PACS Gradient Update
We first formally establish the gradient dynamics of PACS. We demonstrate that minimizing the supervised cross-entropy loss over the score function $\psi(q,o;\pi_{\theta})$ results in a policy update mathematically equivalent to the policy gradient. Specifically, the effective advantage function in this update is dynamically composed of the cross-entropy loss term and the reward prediction error.
Proposition A.1 (Gradient Derivation).
The gradient of the PACS objective function $\mathcal{L}_{\text{PACS}}(\theta)$ with respect to the policy parameters $\theta$ can be formulated as a policy gradient update weighted by an effective advantage function derived from the prediction error and the cross-entropy loss.
Proof.
Consider the PACS objective defined as the expected binary cross-entropy loss:
$\mathcal{L}_{\text{PACS}}(\theta)=-\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(\cdot|q)}\left[\ell_{\text{BCE}}(R(q,o),\sigma(\psi_{\theta}(q,o)))\right]$
[…]
To compute the gradient $\nabla_{\theta}\mathcal{L}_{\text{PACS}}(\theta)$, we apply the log-derivative trick to the expectation term and the chain rule to the integrand:
$\nabla_{\theta}\mathcal{L}_{\text{PACS}}=-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\cdot\ell_{\text{BCE}}(q,o;\pi_{\theta})+\nabla_{\theta}\ell_{\text{BCE}}(q,o;\pi_{\theta})\right]$
For the second term, $\nabla_{\theta}\ell_{\text{BCE}}(q,o;\pi_{\theta})$, we apply the chain rule with respect to the score $\psi_{\theta}$. Noting that the derivative of the sigmoid function is $\sigma^{\prime}=\sigma(1-\sigma)$ and $\frac{\partial\ell_{\text{BCE}}}{\partial\hat{y}}=\frac{\hat{y}-y}{\hat{y}(1-\hat{y})}$, we obtain:
$\nabla_{\theta}\ell_{\text{BCE}}$ $=\frac{\partial\ell_{\text{BCE}}}{\partial\sigma}\frac{\partial\sigma}{\partial\psi}\nabla_{\theta}\psi_{\theta}$
$=\frac{R-\sigma(\psi_{\theta})}{\sigma(\psi_{\theta})(1-\sigma(\psi_{\theta}))}\cdot\sigma(\psi_{\theta})(1-\sigma(\psi_{\theta}))\cdot\nabla_{\theta}\psi_{\theta}$
$=(R(q,o)-\sigma(\psi_{\theta}))\nabla_{\theta}\psi_{\theta}$
Given the definition of $\psi_{\theta}$, its gradient is $\nabla_{\theta}\psi_{\theta}=\beta\nabla_{\theta}\log\pi_{\theta}(o|q)$ (since $\pi_{\text{ref}}$ is fixed). Substituting this back yields:
$\nabla_{\theta}\ell_{\text{BCE}}=\beta(R(q,o)-\sigma(\psi_{\theta}))\nabla_{\theta}\log\pi_{\theta}(o|q)$
Combining both terms, we obtain the final gradient expression:
$\nabla_{\theta}\mathcal{L}_{\text{PACS}}=-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\left[\underbrace{\left(\ell_{\text{BCE}}(q,o;\pi_{\theta})+\beta(R(q,o)-\sigma(\psi_{\theta}))\right)}_{\text{Effective Advantage }A_{\text{PACS}}}\nabla_{\theta}\log\pi_{\theta}(o|q)\right]$
This result confirms that PACS performs a valid policy gradient update. The effective advantage $A_{\text{PACS}}$ naturally incorporates a “critic” signal (the residual $R-\sigma(\psi_{\theta})$) which vanishes as the model’s prediction aligns with the ground truth, providing a variance-reduced update signal compared to standard policy gradients with statistical baselines. ∎
A.2 Implicit Regularization and Robustness to Entropy Collapse
A key empirical finding of this work is that PACS effectively prevents entropy collapse (the degeneration of the policy into a deterministic distribution), a common failure mode in standard RLVR methods like GRPO when explicit KL penalties are omitted. In this section, we provide a theoretical justification for this robustness. We prove that the structural parameterization of PACS imposes an implicit regularization that constrains the optimal policy to a Gibbs distribution, thereby preserving exploration.
Proposition A.2 (Implicit Gibbs Regularization).
Assuming sufficient model capacity, the global minimum of the PACS objective function is achieved if and only if the policy $\pi_{\theta}$ follows a Gibbs distribution anchored to the reference policy $\pi_{\text{ref}}$. This stands in contrast to standard unregularized RL objectives, where the global minimum is a deterministic Dirac distribution.
Proof.
We first analyze the optimality condition for the PACS objective. Since $\mathcal{L}_{\text{PACS}}$ is a strictly convex binary cross-entropy loss with respect to the logits $\psi_{\theta}$, for any query-response pair $(q,o)$, the loss is minimized when the predicted probability matches the expected target reward:
$\sigma(\psi_{\theta}^{*}(q,o))=\mathbb{E}_{\text{data}}[R(q,o)]$
In the context of RLVR with deterministic ground-truth verification, this implies the model attempts to predict the correctness of $o$. Inverting the sigmoid function, the optimal score function $\psi_{\theta}^{*}$ must satisfy:
$\psi_{\theta}^{*}(q,o)=\sigma^{-1}(\mathbb{E}[R(q,o)])$
Crucially, in PACS, the score function is not a free parameter but is structurally coupled to the policy via the definition $\psi_{\theta}(q,o)=\beta(\log\pi_{\theta}(o|q)-\log\pi_{\text{ref}}(o|q))$. Substituting this definition into the optimality condition yields:
$\beta\log\frac{\pi_{\theta}^{*}(o|q)}{\pi_{\text{ref}}(o|q)}=\sigma^{-1}(\mathbb{E}[R(q,o)])$
Exponentiating both sides and rearranging terms, we derive the form of the optimal policy $\pi_{\theta}^{*}$:
$\pi_{\theta}^{*}(o|q)=\pi_{\text{ref}}(o|q)\exp\left(\frac{\sigma^{-1}(\mathbb{E}[R(q,o)])}{\beta}\right)$
Comparison with Standard RL: Consider a standard RL objective $J=\mathbb{E}_{o\sim\pi}[R(o)]$ (often optimized by baselines without explicit KL). The global maximum of $J$ occurs at a deterministic Dirac delta distribution: $\pi_{\text{std}}^{*}=\delta(o-\operatorname*{arg\,max}_{o}R(o))$, which represents total entropy collapse.
In contrast, the optimal policy for PACS derived above is a Gibbs distribution. The term $\pi_{\text{ref}}(o|q)$ acts as a base measure, ensuring that the optimized policy retains the support and distributional characteristics of the pre-trained model. Even without an explicit KL loss term in the training objective, the specific parameterization of $\psi_{\theta}$ enforces this structural constraint. This guarantees that PACS inherently balances reward maximization with adherence to the reference distribution, theoretically explaining the superior diversity and resistance to mode collapse observed in our experiments. ∎
Appendix B Baselines
B.1 Proximal Policy Optimization Algorithms (PPO)
@@
$\mathcal{J}_{\text{PPO}}(\theta)=$ $\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(\cdot|q)}$
$\frac{1}{|o|}\sum_{t=1}^{|o|}\left\{\min\left[\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}A_{t},\text{clip}\left(\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})},1-\varepsilon,1+\varepsilon\right)A_{t}\right]\right\}$
@@
B.2 Group Relative Policy Optimization (GRPO)
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\\ \Bigg\{\min\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}A_{i,t},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)A_{i,t}\right]-\beta D_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\Bigg\}$ (9)
@@
Appendix C Hyperparameters Settings
@@ C.2 Training Hyperparameters
Max prompt length 1024
Max response length 8,192
@@
Micro batch size per GPU 8
@@
Total training steps 120 for Qwen3-4B, 320 for Qwen3-8B
@@ C.3 Inference Hyperparameters
Tensor parallel size 1
@@
Appendix D Detailed Experimental Results
This section contains additional experimental results that are omitted from the main text due to space constraints.
D.1 Results of Qwen3-4B
Detailed performance metrics for Qwen3-4B on the AMC 23, AIME 2024, AIME 2025, and BeyondAIME benchmarks are presented below.
AMC 23 ($\text{pass@}k$) AIME 2024 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
Base 82.30 91.15 93.33 95.34 97.38 99.20 46.48 61.00 66.15 70.21 73.77 77.21
PPO 86.33 93.93 95.48 96.37 97.07 97.47 52.21 70.64 77.03 81.12 82.94 83.32
GRPO 86.95 94.69 95.98 97.08 98.44 99.70 52.84 70.71 76.31 79.97 82.05 83.13
PACS 90.45 96.01 96.96 97.40 97.50 97.50 55.10 72.91 77.93 80.74 82.25 83.13
   - w/o weight 89.18 94.91 95.80 96.54 97.18 97.48 54.56 72.60 78.32 81.68 83.13 83.33
AIME 2025 ($\text{pass@}k$) BeyondAIME ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
Base 34.71 49.27 54.57 58.40 62.41 66.81 17.33 24.92 29.31 34.37 39.91 45.29
PPO 40.34 54.87 60.52 66.36 71.58 75.24 21.82 32.97 39.06 45.46 51.28 56.01
GRPO 40.18 54.87 61.23 67.69 73.48 77.34 24.39 37.90 44.31 49.52 53.79 57.58
PACS 45.63 61.00 66.18 70.14 72.91 74.80 27.16 41.66 47.96 53.41 57.91 62.10
   - w/o weight 44.90 59.10 64.08 68.11 71.37 73.09 25.89 39.11 45.28 51.18 56.53 61.12
@@ D.1 Results of Qwen3-4B
Table 6: Results of Qwen3-4B trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best. Due to space constraints, results on MATH 500 are shown in table 7.
D.2 MATH500
MATH500 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
Base 91.00 95.75 96.86 97.68 98.21 98.50
PPO 93.25 97.06 97.86 98.40 98.80 99.13
GRPO 93.91 97.48 98.20 98.74 99.17 99.43
PACS 94.80 97.87 98.47 98.93 99.30 99.49
   - w/o weight 94.35 97.70 98.41 98.92 99.23 99.35
@@ D.2 MATH500
Table 7: Results of Qwen3-4B on MATH500 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
MATH500 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
Base 89.94 95.20 96.54 97.54 98.27 98.75
PPO 93.78 97.16 97.88 98.38 98.90 99.40
GRPO 95.15 98.08 98.67 99.13 99.50 99.69
PACS 95.09 98.09 98.71 99.16 99.50 99.72
   - w/o weight 95.32 98.06 98.66 99.13 99.46 99.65
@@
Table 8: Results of Qwen3-8B on MATH500 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
We further analyze the performance on the MATH500, as detailed in tables 7 and 8. On the Qwen3-4B scale, PACS demonstrates a consistent and clear advantage over the baselines. Specifically, it achieves a $\text{pass@}1$ score of 94.80%, surpassing GRPO (93.91%) by roughly 0.9 points, and maintains this lead across all $k$ values, reaching 99.49% at $\text{pass@}64$. On the Qwen3-8B scale, performance on this benchmark approaches saturation, with base models already scoring near 90%. Despite this limited room for improvement, PACS remains highly competitive. It significantly outperforms PPO and matches the strong performance of GRPO at $\text{pass@}1$ (95.09% vs. 95.15%), while marginally surpassing it at higher sample counts (e.g., $k\geq 4$). This confirms that PACS retains its stability and effectiveness even on relatively simpler tasks where the solution space is well-explored.
D.3 Ablation of Different Advantage Estimators